a curve is defined parametrically by the equations x=2sintheta y=cos^2(theta) find the equation of the tangent to this curve at (2;0)


1
Expert's answer
2021-01-12T16:08:03-0500

x = "2sin\theta"  and y = "cos^2\theta"   at (2, 0)

The first thing that we should do is find the derivative so we can get the slope of the tangent line.

             "\frac{dy}{dx} = \large\frac{\large\frac{dy}{d\theta}}{\large\frac{dx}{d\theta}}" "= \large\frac{2cos\theta}{-2sin\theta cos\theta}" "=\large -\frac{1}{sin\theta}"

At this point we’ve got a small problem. The derivative is in terms of "\theta" and all we’ve got is an x-y­ coordinate pair. The next step then is to determine the value(s) of "\theta" which will give this  point. We find these by plugging the x and y values into the parametric equations and solving for "\theta"

           2 = 2"sin\theta"    "\to"    "\theta = \large\frac{\pi}{2} + 2\pi k"

           0 = "cos^2\theta"     "\to"    "\theta" "= \large\frac{\pi}{2} + \pi k"


Since we already know the x and y coordinates of the point all that we need to do is find the slope of the tangent line.

             "m = \large\frac{dy}{dx} \large|_{x = -\large\frac{\pi}{2}} =" 1


"y - y_0 = m (x - x_o)"

y - 0 = 1* (x - 2)

y = x - 2